# Ultrafast optical excitation of magnons in 2D antiferromagnetic semiconductors via spin torque mediated by unbound electron-hole pairs and excitons: Signatures in magnonic charge pumping

### v4

@@ III Conclusions and Outlook
Using time-dependent quantum transport [75, 76] combined with the classical LLG equation, [68, 69] as well as extended [19] to include Coulomb-binding of photoexcited electrons and holes into excitons, we explain that the microscopic mechanism behind recent experimental observations of optical excitation of coherent [38, 39] magnons in 2D AF semiconductors [3, 4, 5, 14, 11, 9, 6] or their vdW heterostructures [14] is a nonequilibrium spintronic effect of spin torque type [83, 122]. The strongest experimental support for our theory comes from the fact that without applying a magnetic field to induce sufficient noncollinearity [i.e., their canting in Fig. 1] between LMMs of two layers of CrSBr, no magnon excitation is detected in Refs. [3, 4, 5]. This is because for collinear or nearly collinear (such as due to thermal fluctuations) LMMs, STT in Eq. (1) is vanishingly small. Thus, our microscopic theory displaces the need for a phenomenological “impulsive perturbation” [5, 14] typically invoked within the classical LLG equation alone (i.e., without considering photoexcited electrons), as a standard tool employed to interpret recent experiments. It also connects ultrafast carrier dynamics in 2D magnetic materials to conventional spintronics, where STT is a cornerstone effect but is usually studied [83, 122, 84, 85] in magnetic heterostructures driven by a small DC bias voltage. Finally, it predicts additional charge pumping by excited magnons (also discussed in conventional spintronics literature [54, 52, 55, 53] but with GHz fields driving magnons and, therefore, of much lower frequency than fsLP), as well as signatures of excitons in high harmonics [93] [Figs. 4 and  5] of pumped charge current. Thus, such unexplored signals of excited magnons and their interaction with excitons could be exploited in future experiments.
@@
Another recently developed quantum-classical formalism [65]—where the Schrödinger equation, employing the first-principles derived effective Hamiltonian in terms of composite bosons describing excitons, is coupled self-consistently to the classical LLG equation for LMMs—has found excitation of magnons and total magnetization reversal in ultrafast light-driven CrI3. The same approach could, in principle, be applied to other 2D AF semiconductors. However, it does not take into account unbound electron-hole pairs, which coexist [20] with excitons when using above-gap light. For example, while experiments on NiPS3 have observed magnon excitation for subgap light tuned to exciton energy and, therefore, mediated by it, in the case of CrSBr, no magnons are detected for subgap light [123]. This suggests that considering both unbound electron-hole pairs and excitons is required to understand above-gap light excitation of CrSBr [20], as naturally captured [16, 17, 19] by TDNEGF formalism but absent in wavefunction-based approaches [65]. Our TDNEGF+LLG+EX formalism [Fig. 2] formulated for open quantum systems [such as the one in Fig. 1] also naturally computes spin and charge currents [Fig. 5] pumped out of the system or STT within the system [see movie provided as the SM].
@@
Specifically for CrSBr, the very recent experiments have observed additional effects associated with exciton-magnon coupling, not found in other 2D AF semiconductors [6], where either magnons modulate exciton energy (by a few meV) [3]; or HHG [5] of magnons are generated at frequencies $n\omega_{b}$, with $n$ reaching surprisingly large $n\gtrsim 20$. Regarding the former, our self-consistent loop in Fig. 2, where magnon presence adds a time-dependent term into the electronic Hamiltonian, could explore this effect by focusing on exciton properties, rather than on magnon properties, which are the focus of the present study. Regarding the latter, our theory reproduces only “bright magnon” at frequency $\omega_{b}$, while not capturing HHG [Fig. 3]. Since the LLG equation is nonlinear, it can describe [71, 35] magnon-magnon interactions as one of the key ingredients [124] for large $n$. Such interactions are enhanced [34] by experimentally induced noncollinearity (i.e., canting in Fig. 1) of LMMs within CrSBr [5]. Although Fig. 3 suggest that LLG description is insufficient to capture the HHG of magnons in CrSBr, we offer a hint [Fig. 6(c)] in the SM toward possible ingredients required to describe this effect—there we find that increasing $J_{K}$ to an unrealistically large value (i.e., ten times larger than used in Figs. 35, as a parameter whose precise value is difficult to extract from first-principles calculations [65]) leads to HHG but at noninteger $n$. We anticipate that adding quantum corrections [125, 126] to the LLG equation and/or by treating excitons beyond tMFT could explain the experimentally observed [5] HHG of magnons and the role of many-body interactions [127] with excitons in this process. For example, within TDNEGF formalism, we can include additional self-energies to go beyond the tMFT description of excitons, as well as systematically improve them [15, 77, 16, 17]. We relegate such extensions of the TDNEGF+LLG+EX framework to future studies.
@@ Acknowledgements.
Supplemental Material
Besides movie file—animating dynamics of LMMs  [^1] Throughout this Supplemental Material, we use the same acronyms as in the main text: 1D-one-dimensional; 2D-two-dimensional; AF-antiferromagnetic; CAR-central active region; EM-electromagnetic; FF-far field; FFT-fast Fourier transform; FM-ferromagnetic; fsLP-femtosecond laser pulse; HHG-high harmonic generation; $L$-left; LMM-localized magnetic moment; NN-nearest-neighbor; NM-normal metal; $R$-right; STT-spin-transfer torque; TB-tight-binding; TDNEGF+LLG-time-dependent nonequilibrium Green’s functions + Landau-Lifshitz-Gilbert; TDNEGF+LLG+EX-same as TDNEGF+LLG but with inclusion of excitons; tMFT-time-dependent mean field theory. and STT exerted on them, as well as charge currents generated by dynamical LMMs (i.e., magnons they comprise)— below we include details of our quantum and classical Hamiltonian models and methods for computing EM radiation via the Jefimenko equations, as well as one additional Fig. 6.
Appendix A Additional Figure
@@
One additional Fig. 6 of this SM consists of three panels. Its first panel, Fig. 6(a), analyzes the effect of photoexcited electrons—for simplicity, without their binding into excitons [$U=0$ in Eq. (14)]—on the time evolution of optically excited magnons. In particular, the goal is to understand the origin of magnon peaks in Figs. 3(b) and  3(d) of the main text below the frequency $\omega_{b}$ of “bright magno.” For this purpose, we allow photoexcited electrons to ignite magnons microscopically (instead of adding phenomenological “impulsive perturbation” [5, 14] into the LLG equation) via STT exerted by electronic spin current [83, 84, 85], but then we immediately switch off Kondo exchange interaction between flowing electronic spins and LMMs. In other words, we use the full self-consistent loop [Fig. 2 of the main text] in the TDNEGF+LLG framework before $t=15$ ps in Fig. 6(a), and then for $t\geq 15$ ps $J_{K}(t)=0$, so that the TDNEGF and LLG parts of the loop are disconnected. Such calculations lead to a longer lifetime of magnon at frequency $\omega_{b}$ [5] and its second harmonic $2\omega_{b}$ in Fig. 6(a)—compare the length of bright lines in Fig. 6(a) at these two frequencies with their counterparts from Fig. 3(b) of the main text. The longer lifetime is the consequence of removing nonlocal damping [79, 110, 111, 112, 113, 114, 115] by $J_{K}=0$ for $t\geq 15$ ps, which is otherwise provided by the fermionic bath of electrons within TDNEGF calculations of the TDNEGF+LLG loop. In contrast, such damping is present in all figures of the main text during the whole evolution time. Note that the counterpart of the magnon at frequency $2\omega_{b}$, as clearly visible in Fig. 6(a), is of vanishingly short lifetime in Fig. 3(b) in the main text due to nonlocal damping being operative in the latter case. The remaining finite lifetime of the $2\omega_{b}$ magnon in Fig. 6(a) is due to retained conventional local Gilbert damping of strength $\alpha_{G}$ [Eq. (3) of the main text]. Importantly, switching off Kondo exchange interaction removes [Fig. 6(a)] bright lines of excited magnons that are otherwise present in Fig. 3(b) of the main text at frequencies below $\omega_{b}$. This finding confirms that magnon peaks below $\omega_{b}$ are due [58] to photoexcited electrons, which are explicitly included in our TDNEGF+LLG framework but are absent from the LLG equation used on its own in Fig. 6(a) for $t\geq 15$ ps.
Figure 6: Power spectrum of windowed FFT [90, 91], $|N^{z}(t_{c},\omega)|^{2}$, of the $z$-component of the Néel vector [Eq. (8) in the main text] for: (a) and (c) TDNEGF+LLG-computed dynamics in the absence of excitons ($U=0$); or (b) the same dynamics probed experimentally in CrSBr at externally applied magnetic field $B=0.3$ T [note that the same data from panel (b) was analyzed in Fig. 4b of experimental Ref. [5] but using standard FFT]. The dynamics of the Néel vector in panel (a) is initiated by photoexcited electrons, i.e., in the same fashion as in Fig. 3 of the main text, but then for $t\geq 15$ ps we switch off ($J_{K}=0$) Kondo exchange interaction between flowing electronic spins and LMMs. In panel (c), we increase the value of $J_{K}$ used in the main text by ten times. Note that $\omega_{b}$ is larger in our calculations within panels (a) and (c) than in the experimental data of panel (b), so for easy comparison we rescaled the ordinate of panel (b).
@@
For comparison, we apply windowed FFT [90, 91] to experimental data from Ref. [5] generated in the course of detection of optically excited magnons in CrSBr. Such analysis replicates in Fig. 6(b) “bright magnon” at frequency $\omega_{b}$, while additionally (in comparison with standard FFT performed in Ref. [5]) revealing its long lifetime. We also unravel shorter-living magnons at high harmonics $n\omega_{b}$. However, the power spectrum of windowed FFT is too noisy in Fig. 6(b) for frequencies below $\omega_{b}$ to allow us to extract additional magnon peaks expected [Figs. 3(b) and 3(d) of the main text] to be excited due to flowing electronic spins interacting with LMMs. We speculate that magnon peaks below $\omega_{b}$ frequency could be present in experiments, but they are short-lived and, therefore, invisible to the standard FFT employed in Ref. [5].
@@
Finally, in contrast to Fig. 6(a), where $J_{K}$ is abruptly reduced to zero for $t\geq 15$ ps, in Fig. 6(c) we keep it nonzero for all times while increasing its value ten times ($J_{K}=0.1$ eV) when compared to the same Kondo exchange interaction between flowing electronic spins and LMMs used in the main text ($J_{K}=0.01$ eV). This choice leads to HHG in Fig. 6(c), but at frequencies that are not exact integer multiples $n\omega_{b}$. While this finding does not match integer HHG observed experimentally [5] and reproduced in Fig. 6(b), it does offer a hint for future extensions of the TDNEGF+LLG+EX framework and its application to HHG of magnons, as discussed in the main text.
Appendix B Models and Methods
Our TDNEGF+LLG+EX calculations are fully microscopic (i.e., Hamiltonian-based), requiring only two Hamiltonians as an input. These two Hamiltonians describe bare degrees of freedom—quantum Hamiltonian, $\hat{H}(t)$, for electrons; and the classical one, $\mathcal{H}(t)$, for LMMs. In this study, we focus on essential features of bilayer 2D AF semiconductors [12, 11, 56] by constructing a model quantum Hamiltonian [Eq. (11)] that captures two monolayers, a semiconducting bandgap for each of them, ferromagnetic (FM) intralayer and AF interlayer ordering of LMMs, and possible Coulomb interaction effects [15, 19] binding conduction-band electrons and valence-band holes into excitons. Accordingly, our simulation setup in Fig. 1 of the main text consists of two 1D TB chains [46, 47, 51] with two electronic orbitals per site $i$, as well as one LMM per site $i$ described by the unit vector $\mathbf{M}_{i}(t)$. Our usage of 1D chains is inspired by CrSBr [46, 47, 48, 49, 50, 51] as a highly anisotropic 2D magnetic material [1, 2] formed [46, 47, 51] by 1D atomic chains with interlayer AF ordering (A-type) and intralayer FM ordering. The Curie temperature of FM ordering within a single monolayer is $\simeq 150$ K, while AF ordering of LMMs between adjacent layers has a Néel temperature of $\simeq 130$[49]. Note that angle-resolved photoemission spectroscopy reveals the complex electronic band structure of CrSBr, whose accurate theoretical description requires first-principles methods, such as DFT+U [51] or self-consistent GW methodology [49, 10, 23].
The classical Hamiltonian, entering into the LLG Eq. (3) in the main text, is given by
Such an effective classical Heisenberg Hamiltonian of LMMs can be extracted from first-principles calculations [129, 66, 130] or by fitting magnon spectra from neutron scattering data [48]. Here $J^{\rm AF}=0.0195\ \mathrm{eV}$ and $J^{\rm FM}=0.15\ \mathrm{eV}$ are AF and FM exchange couplings, respectively; $K_{x}=0.021\ \mathrm{eV}$ and $K_{z}=0.057\ \mathrm{eV}$ specify the magnetic anisotropy along the $x$- and $z$-axis, respectively; $J_{K}=0.01\ \mathrm{eV}$ is Kondo exchange interaction between spins of flowing electrons and LMMs [86, 87]; $i\in 1$ ($i\in 2$) means that the site $i$ belongs to the first (second) layer in Fig. 1 of the main text; and $\langle\ldots\rangle$ denotes NN sites. Note that $J^{\rm AF},\ J^{\rm FM},\ K_{x},\ K_{z}$ parameters were also employed to interpret experiments [3, 4, 5], but we rescale their values by a multiplicative factor $\times 10^{3}$ to make the total TDNEGF+LLG+EX simulation time of $\sim 10$ ps duration. Nevertheless, even with this adjustment, the characteristic energy scale of the AF background remains below electron kinetic energy, $\gamma_{0}/J^{\rm{AF}}\simeq 10^{2}$.
Here $\hat{H}_{\mathrm{intra}}$ describes 1D TB chains within layer 1 or 2 in Fig. 1 in the main text
$\hat{H}_{\mathrm{intra}}=\Delta/2\sum_{i\in 1,2}(\hat{c}_{i}^{c\dagger}\hat{c}^{c}_{i}-\hat{c}_{i}^{v\dagger}\hat{c}^{v}_{i})$
where indices $c$ and $v$ stand for orbitals that give rise to the conduction and valence bands of each chain; $\hat{c}_{i}^{a\dagger}=(\hat{c}_{i\uparrow}^{a\dagger}\ \ \hat{c}_{i\downarrow}^{a\dagger})$ is the row vector containing operators $\hat{c}_{i\sigma}^{a\dagger}$ creating electrons with spin $\sigma=\uparrow,\downarrow$ in orbital $a=c,v$ hosted by site $i$; $\Delta$ is the onsite potential opening bandgap between the two bands $\Delta=3\ \rm{eV}$; and $\gamma_{0}=1\ \mathrm{eV}$ is the hopping parameter between the NN sites. The spin density operator in Eq. (11) is given by $\mathbf{\hat{s}}_{ia}=\hat{c}_{i}^{a\dagger}\hat{\bm{\sigma}}\hat{c}^{a}_{i}$ for $a=c,v$. The fsLP is introduced in Eq. (B) via the Peierls phase [131, 132], $\chi_{ij}(t)=z_{\max}\exp[-\left(t-t_{p}\right)^{2}/(2\sigma_{\text{light }}^{2})]\sin(\omega_{0}t)$, where the electric field of the pulse is $\mathbf{E}(t)=-\partial_{t}\mathbf{A}(t)$, $z_{\max}=eaA_{\rm max}/\hbar=0.1$ is a dimensionless parameter quantifying the intensity of the pulse using $A_{\rm max}$ as the amplitude of the vector potential, $\omega_{0}$ is the central frequency of fsLP, and $a$ is the lattice spacing. Besides the Peierls phase, fsLP is additionally [15] introduced via hopping $\gamma_{\mathrm{P}}(t)=\mathbf{d}\cdot\mathbf{E}(t)$ in Eq. (B). This term describes those interband transitions that are driven by the dipole interaction with the electric field [15], where $\mathbf{d}$ is the expectation value of the dipole operator, $\mathbf{d}=e\langle i,c|\mathbf{\hat{r}}|i,v\rangle$. The term $\hat{H}_{\mathrm{inter}}$ is given by
and it describes hopping with parameters $\gamma_{c}=0.5\gamma_{0}$ or $\gamma_{v}=0.5\gamma_{0}$ between $c$ or $v$ orbitals, respectively, located at NN sites of two different chains. Finally, the inter-orbital Coulomb interaction [15, 19]
describes how two electrons on two different orbitals at the same site $i$ interact with each other. The same Hubbard-type term was employed in prior studies [15, 19] to describe exciton formation on TB lattice. We decouple it via tMFT [19] (otherwise, its beyond-tMFT treatment requires computationally much more expensive evaluation of Feynman diagrams on the Keldysh contour [15]) as follows:
Here the first two terms correspond to Hartree and the latter two to Fock approximation, where the order parameter of the excitonic condensate [36] is given by $\phi_{i}(t)=\langle\hat{c}_{i}^{v\dagger}(t)\hat{c}^{c}_{i}(t)\rangle$. This procedure requires us to self-consistently compute $\phi_{i}(t)$, which we obtain from the off-diagonal elements of ${\bm{\rho}}^{\rm neq}(t)$ as a part of TDNEGF calculations within the self-consistent loop illustrated in Fig. 2 of the main text. We note that in this approach we do not rely on the effective excitonic Hamiltonians written in terms of operators of composite bosons [65], which can be cumbersome to derive [133]. Instead, we use “bare” degrees of freedom, so that both unbound electron-hole pairs and their binding into excitons by the Coulomb interaction are considered through the electronic density matrix, as naturally included [19, 15, 16, 17] when using the nonequilibrium density matrix and/or TDNEGF formalism.
The AF semiconductor CAR in Fig. 1 of the main text is attached to $L$ and $R$ NM leads modeled as semi-infinite ideal 1D TB chains with one orbital per site. The chemical potential of the macroscopic reservoirs into which NM leads terminate is identical (i.e., no bias voltage is applied) and chosen as $\mu_{L}=\mu_{R}=E_{F}=0$.
Note that Jefimenko Eq. (16) can be viewed [102] as proper (i.e., time-retarded) time-dependent generalizations of the Coulomb law. Here, $t_{r}\equiv t-|\mathbf{r}-\mathbf{l}|/c$ emphasizes retardation in the response time due to relativistic causality [101, 103]. Additionally, we adapt [104, 73] Eq. (16) to utilize time-dependent bond [116] charge currents as the source of EM radiation, $I_{ia\rightarrow jb}(t)$ [Eq. (17)]. They are the counterpart (on the TB lattice) of local current density in continuous space. The bond currents $I_{ia\rightarrow jb}$ are assumed to be spatially homogeneous along the path $P_{ia\rightarrow jb}$ from orbital $a$ at site $i$ to orbital $b$ at site $j$ [116, 78, 104], which is composed of a set of points $l\in P_{ia\rightarrow jb}$. Here $N_{B}$ is the number of bonds $ia\rightarrow jb$, and since we use $N=10$ black and gray sites [Fig. 1 of the main text] in each layer of CrSBr, $N_{B}=36$ in our calculations. We obtain bond charge currents as

### v5

@@ IV.1 Conclusions
Using time-dependent quantum transport [56, 57] combined [62, 61] with the classical LLG equation [59, 60], as well as extended [19, 15] to include Coulomb-binding of photoexcited electrons and holes into excitons, we explain that the microscopic mechanism behind recent experimental observations of optical excitation of coherent [38, 39] magnons in 2D AF semiconductors [3, 4, 5, 14, 11, 9, 6] or their vdW heterostructures [14] is a nonequilibrium spintronic effect of spin torque type [132, 133]. The strongest experimental support for our theory comes from the fact that without applying a magnetic field to induce sufficient noncollinearity [i.e., their canting in Fig. 1] between LMMs of two layers of CrSBr, no magnon excitation is detected in Refs. [3, 4, 5]. This is because for collinear or nearly collinear (such as due to thermal fluctuations [80, 91]) LMMs, STT in Eq. (1) is vanishingly small. Thus, our microscopic theory displaces the need for a phenomenological “impulsive perturbation” [5, 14] typically invoked within the classical LLG equation alone (i.e., without considering photoexcited electrons), as a standard tool employed to interpret recent experiments. It also connects ultrafast carrier dynamics in 2D magnetic materials to conventional spintronics, where STT is a cornerstone effect but is usually studied [132, 133, 85, 86] in magnetic heterostructures driven by a small DC bias voltage or current pulses. Finally, it predicts additional charge pumping by excited magnons (discussed in prior spintronic literature [54, 52, 55, 53], but for magnetic systems driven by GHz fields rather than light), as well as signatures of excitons in high harmonics [106] [Figs. 4 and  5] of pumped charge current. Thus, such unexplored signals of excited magnons and their interaction with excitons could be exploited in future experiments. Regarding which signal to choose for future experiments, both detection of pumped charge current in electrical contacts or its EM radiation contain nearly identical information [compare Fig. 4 vs.  5]. So, the choice will largely depend on how easy it is to fabricate an optoelectronic device where 2D material is attached to electrical contacts. If that is possible (see, e.g., device in Fig. 1 of Ref. [134]), and pumped current has spectral content in the GHz or smaller frequency range; we recommend analyzing such current flowing into an external attached circuit. If not, one can detect far-field EM radiation in the GHz or THz range of frequencies, which is a well-established experimental technique [128, 129, 131, 130] in ultrafast spintronics when detection of excited spin and charge currents via an external circuit is cumbersome or impossible.
IV.2 Outlook
@@
Another recently developed quantum-classical formalism [73]—where the Schrödinger equation, employing the first-principles derived effective Hamiltonian in terms of composite bosons describing excitons, is coupled self-consistently to the classical LLG equation for LMMs—has found excitation of magnons and total magnetization reversal in ultrafast light-driven CrI3. The same approach could, in principle, be applied to other 2D AF semiconductors. However, it does not consider unbound electron-hole pairs, which coexist [20] with excitons when using above-gap light. For example, while experiments on NiPS3 have observed magnon excitation for subgap light tuned to exciton energy and, therefore, mediated by it, in the case of CrSBr, no magnons are detected for subgap light [135]. This suggests that considering both unbound electron-hole pairs and excitons is required to understand above-gap light excitation of CrSBr [20], as naturally captured [16, 17, 19] by the TDNEGF formalism but absent in wavefunction-based approach of Ref. [73]. Our TDNEGF+LLG+EX formalism [Fig. 2] formulated for open quantum systems [such as the one in Fig. 1] also naturally computes spin and charge currents [Fig. 5] pumped out of the system or STT within the system (see movie provided as the SM 11footnotemark: 1), which cannot be studied by time-evolving wavefunction of closed quantum systems.
@@
Specifically for CrSBr, the very recent experiments have also observed additional effects associated with exciton-magnon coupling, not found in other 2D AF semiconductors [6], where either magnons modulate exciton energy (by a few meV) [3]; or HHG [5] of magnons are generated at frequencies $n\omega_{b}$, with $n$ reaching surprisingly large values $n\gtrsim 20$. Regarding the former, our self-consistent loop in Fig. 2, where magnon presence adds a time-dependent term into the electronic Hamiltonian, could explore this effect by focusing on exciton properties, rather than on magnon properties in the focus of the present study. Regarding the latter, our theory reproduces only “bright magnon” at frequency $\omega_{b}$, while not capturing [Fig. 3] HHG of magnons at $n\omega_{b}$ frequencies. Since the LLG equation is nonlinear, it can in principle describe [77, 35] magnon-magnon interactions as one of the key ingredients [136] for large $n$. Such interactions are enhanced [34] by experimentally induced noncollinearity (i.e., canting in Fig. 1) of LMMs within CrSBr [5]. Although Fig. 3 suggest that LLG description is insufficient to capture the HHG of magnons in CrSBr, we offer a hint in the Appendix [Fig. 6(c)] toward possible ingredients required to describe this effect—there we find that increasing $J_{K}$ to an unrealistically large value (i.e., ten times larger than used in Figs. 35, as a parameter whose precise value is difficult to extract from first-principles calculations [73]) does yield HHG, but at noninteger values $n$.
Figure 6: Power spectrum of windowed FFT [116, 117], $|N^{z}(t_{c},\omega)|^{2}$, of the $z$-component of the Néel vector [Eq. (17)] for: (a) and (c) TDNEGF+LLG-computed dynamics in the absence of excitons ($U=0$); or (b) the same dynamics probed experimentally in CrSBr at externally applied magnetic field $B=0.3$ T [note that the same data from panel (b) was analyzed in Fig. 4b of experimental Ref. [5] but using standard FFT]. The dynamics of the Néel vector in panel (a) is initiated by photoexcited electrons, i.e., in the same fashion as in Fig. 3, but then for $t\geq 15$ ps we switch off ($J_{K}=0$) Kondo exchange interaction between flowing electronic spins and LMMs. In panel (c), we increase the value of $J_{K}$ used in Fig. 3 by ten times. Note that $\omega_{b}$ is larger in our calculations within panels (a) and (c) than in the experimental data of panel (b), so for easy comparison we rescaled the ordinate of panel (b).
We anticipate that replacing the LLG equation with semiclassical theory but capable of capturing nonlinear dynamics that includes nonclassical effects, as proposed very recently in Ref. [79] and applied to antiferromagnets, and/or by treating excitons beyond tMFT could explain the experimentally observed [5] HHG of magnons and the role of many-body interaction [137] with excitons in this process. For example, within the TDNEGF formalism, we can include additional self-energies to surpass the tMFT description of excitons, as well as systematically improve them [15, 58, 16, 17]. Note that tMFT, or different versions of diagrammatic approximations for the self-energy in TDNEGF formalism, have been compared with the exact benchmark obtained from tensor network-based algorithms in Ref. [15], finding that the most important difference between the tMFT and the rest is the lack of proper damping of the induced coherent oscillations. We relegate possible extensions (as delineated in this paragraph) of the TDNEGF+LLG+EX framework to future studies.
@@ Acknowledgements.
Appendix A Additional examples of windowed FFT related to Fig. 3
@@
An additional Fig. 6 applies windowed FFT to two cases with artificially switched off or increased $J_{K}$ parameter, as well as to experimental data of Ref. [5]. This then helps to clarify our results in Fig. 3. Figure 6(a) analyzes the effect of photoexcited electrons—for simplicity, without their binding into excitons [$U=0$ in Eq. (9)]—on the time evolution of optically excited magnons. In particular, the goal is to understand the origin of magnon peaks in Figs. 3(b) and  3(d) below the frequency $\omega_{b}$ of “bright magnon.” For this purpose, we allow photoexcited electrons to ignite magnons microscopically (instead of adding phenomenological “impulsive perturbation” [5, 14] into the LLG equation) via STT exerted by electronic spin current [132, 85, 86], but then we immediately switch off Kondo exchange $J_{K}$ interaction between flowing electronic spins and LMMs. In other words, we use the full self-consistent loop [Fig. 2] in the TDNEGF+LLG framework before $t=15$ ps in Fig. 6(a), and then for $t\geq 15$ ps $J_{K}(t)=0$, so that the TDNEGF and LLG parts of the loop are disconnected. Such calculations lead to a longer lifetime of magnon at frequency $\omega_{b}$ [5] and its second harmonic $2\omega_{b}$ in Fig. 6(a)—compare the length of bright lines in Fig. 6(a) at these two frequencies with their counterparts from Fig. 3(b). The longer lifetime is the consequence of removing nonlocal damping [62, 121, 122, 123, 124, 125, 126] by $J_{K}=0$ for $t\geq 15$ ps, which is otherwise provided by the fermionic bath of electrons within TDNEGF calculations of the TDNEGF+LLG loop. In contrast, such damping is present in Fig. 3 during the whole evolution time. Note that the counterpart of the magnon at frequency $2\omega_{b}$, as clearly visible in Fig. 6(a), is of vanishingly short lifetime in Fig. 3(b) due to nonlocal damping being operative in the latter case. The remaining finite lifetime of the $2\omega_{b}$ magnon in Fig. 6(a) is due to retained conventional local Gilbert damping of strength $\alpha_{G}$ [Eq. (4)]. Importantly, switching off the Kondo exchange interaction removes [Fig. 6(a)] bright lines of excited magnons that are otherwise present in Fig. 3(b) at frequencies below $\omega_{b}$. This finding confirms that magnon peaks below $\omega_{b}$ are due [65] to photoexcited electrons, which are explicitly included in our TDNEGF+LLG framework but are absent from the LLG equation used on its own in Fig. 6(a) for $t\geq 15$ ps.
@@
For comparison, we apply windowed FFT [116, 117] to experimental data from Ref. [5] generated in the course of the detection of optically excited magnons in CrSBr. Such analysis replicates in Fig. 6(b) “bright magnon” at frequency $\omega_{b}$, while additionally (in comparison with standard FFT performed in Ref. [5]) revealing its long lifetime. We also unravel shorter-lived magnons at high harmonics $n\omega_{b}$. However, the power spectrum of windowed FFT is too noisy in Fig. 6(b) for frequencies below $\omega_{b}$ to allow us to extract additional magnon peaks expected [Figs. 3(b) and  3(d)] to be excited due to flowing electronic spins interacting with LMMs. We speculate that magnon peaks below $\omega_{b}$ frequency could be present in experiments, but they are short-lived and, therefore, invisible to the standard FFT employed in Ref. [5].
@@
Finally, in contrast to Fig. 6(a), where $J_{K}$ is abruptly reduced to zero for $t\geq 15$ ps, in Fig. 6(c) we keep it nonzero for all times while increasing its value ten times ($J_{K}=0.1$ eV) when compared to the same Kondo exchange interaction between flowing electronic spins and LMMs used in Fig. 3. This intervention yields HHG in Fig. 6(c), but at frequencies that are not exact integer multiples $n\omega_{b}$. While such a finding does not match integer HHG observed experimentally [5] and reproduced in Fig. 6(b), it does offer a hint for future extensions of the TDNEGF+LLG+EX framework, as discussed in Sec. IV.